OpenStudy (anonymous):

solve -3 + 2x = -9 will give medal

OpenStudy (texaschic101):

-3 + 2x = -9 --- add 3 to both sides 2x = -9 + 3 2x = - 6 x = -6/2 x = -3 check.. -3 + 2(-3) = -9 -3 - 6 = -9 -9 = -9 (correct) x = -3

OpenStudy (uri):

-3 + 2x = -9 2x=-9+3 2x=-6 x=-3 woott!

OpenStudy (muzzack):

do not give answers out!! @nuviaa0203 @Hero and @amistre64

OpenStudy (muzzack):

well what @texaschic101 was right

OpenStudy (anonymous):

can you guys help with one more? @texaschic101

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 } y - 4 = 7\]

OpenStudy (anonymous):

@uri @texaschic101